# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### v1

@@ Abstract
We develop a continuum framework for confining planar QCD ($N_{c}\to\infty$) by quantizing the Fermi-string (1981) degrees of freedom on the rigid Hodge-dual minimal surface constructed in Geometric QCD I. The internal Majorana fermions on the worldsheet provide the algebraic mechanism required by the Makeenko–Migdal loop equations: by the Pauli principle, they enforce cancellations of non-planar self-intersection contributions, leaving the planar factorization structure. The Liouville instability of the original random-surface formulation is avoided because the bulk geometry is rigid and fixed holographically by the boundary loop, with no summation over fluctuating worldsheet metrics.
A key outcome is a formulation in momentum loop space, where coordinate-space cusp/contact singularities are integrated out, and the relevant quark-loop amplitudes admit a finite local limit. A central technical step is an explicit change of variables in the quark-loop phase-space path integral: after gauge fixing the reparametrization symmetry (Virasoro constraint), the reduced loop-space measure is parametrized by boundary twistors and an induced Jacobian, yielding a concrete twistor-string functional integral derived directly from the planar QCD amplitude rather than postulated as an abstract string model.
In the local limit, the theory becomes a confining (analytic) twistor string: a boundary sigma model $S^{1}\to(S^{3}\times S^{3})/S^{1}$ coupled to a holographically determined Liouville field $\rho(z,\bar{z})=\log\!\bigl(|\lambda(z)||\mu(z)|\bigr)$ obtained by analytic continuation of the boundary twistor data (Hilbert/Plemelj reconstruction).
The meson spectrum is encoded in a one-dimensional functional integral over the boundary twistor trajectory $\lambda(z),\mu(z),|z|=1$. While we evaluate this integral numerically using a reweighted Complex Langevin approach—successfully extracting approximately linear Regge trajectories as a proof of concept—our ultimate finding is analytical. By analyzing the monodromies of complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles inside the unit circle. The simplest sector with just one pole has the spectrum $m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$.
@@ 1 Introduction
The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the loop equations. At leading order in the $1/N_{c}$ expansion, the Makeenko–Migdal hierarchy exhibits the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
@@
Remark 1.1 (Equation solving vs. model building).
The starting point of this work is not a proposed “string model of QCD,” but the planar Makeenko–Migdal loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD. The aim is constructive: to exhibit explicit variables and an explicit functional measure in which the loop equations admit a finite continuum solution and yield physical observables. The principal ingredients of the present construction are not adjustable assumptions. The rigid Hodge-dual minimal surface enters because strict additivity and an (anti)self-dual area derivative are algebraically required for compatibility with the loop hierarchy, while the Majorana determinant enters because its Pauli-principle sign structure enforces planar factorization.
@@
The culminating result of this paper is the exact reduction of the meson spectroscopy problem to a one-dimensional functional integral over boundary twistor data $(\lambda(z),\mu(z),\,|z|=1)$. While we initially evaluate this integral numerically via a reweighted Complex Langevin (CLE) approach—successfully extracting approximately linear Regge trajectories as a computational proof of concept—our ultimate finding is analytical. By evaluating the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory. The path integral exactly localizes at critical resonance energies where the Lefschetz thimbles become degenerate. Because the integration variables are coupled holomorphic twistors, this degeneracy intrinsically possesses a corank of 2, mathematically upgrading generic semiclassical branch cuts into the exact simple poles of hadronic states. Planar QCD thus ceases to be a theory of stochastic quantum fluctuations; the extraction of the exact spectrum reduces to a deterministic, classical generalized eigenvalue problem. While we initially evaluate the one-dimensional functional integral over the boundary twistor trajectory numerically using a reweighted Complex Langevin approach—successfully extracting approximately linear Regge trajectories as a proof of concept—our ultimate finding in this paper is analytical.
By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory. The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral. We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent. Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle. This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$. By analytically continuing the action to the boundary and summing over the exact monodromy phases, we obtain an exact Bohr-Sommerfeld quantization condition. For the simplest sector with just one pole, this yields the spectrum:
$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (1.1)
Remarkably, the universally correct Lüscher open-string intercept (+1/12+1/12) emerges directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations.
@@
To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with four specific navigation points that address anticipated physical and computational questions.
@@
Sections 10 to 12 introduce momentum loop space and the twistor string framework. To clarify (ii) the status of Wilson-loop “renormalization,” Sec. 10.1 sharply outlines our perspective: $W[C]$ is kept regularized at a finite Elf mass, physical observables are constructed natively in momentum loop space $W[P]$ (where coordinate-space singularities are integrated out), and only then is the local limit taken. Following this, Sections 11 to  12 execute the central change of variables. Addressing (iii) the parametrization of the quark loop in phase space by boundary values of holomorphic twistors, Sec. 11 details how the linear phase-space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
Section 13 discusses the semiclassical (WKB) limit and the emergence of linear Regge behavior. Section 14.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
Then, we address (iv) the discrete approximation and numerical implementation of the path integral. Section 15 summarizes the CLE evaluation, while E provides exhaustive details on the CLE implementation and topological analysis. Section 16 details the unsupervised phase-information clustering used to isolate the spectral branches.
Finally, we highlight (v) the exact analytic extraction of the mass spectrum. Section 17.1 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory. This last observation reveals the true geometric nature of Planar QCD: the main theoretical result of this work. Section 17.6 discusses the geometric origin of QCD, and Section 18 concludes. The remaining appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C.
@@ 2.3 The Physical Vacuum and Multi-Instanton Resummation
Before we formulate the dynamics of the Fermi string on the minimal surface, we must clarify the physical origin of this rigid geometric background. The Makeenko-Migdal (MM) loop equations serve as the exact quantum equations of motion for Planar QCD. Like any equations of motion, they possess an inherent ambiguity corresponding to the choice of the physical vacuum. The mathematical freedom to multiply the perturbative loop functional by a zero mode exactly reflects this physical vacuum ambiguity.
@@ 3.3 The Virasoro Constraint and Uniformization
In the presence of the Virasoro constraint (3.21), the problem becomes a challenging one and has been investigated over the last two centuries by great mathematicians, starting with Riemann and Hilbert. We give a brief summary of modern state of this theory in Appendix B.
@@ 6.6 Loop equation for fermion determinants
This is clear from the original representation of $\Delta$ as an exponential of the sum over configurations of independent loops. This planar hierarchical set of loops is shown in Fig. 4 .
Figure 4: The hierarchical set of loops, some touching, but never intersecting each other nor self-intersecting
@@ 10.4 Equivalence of loop diffusion operator to the third Magnus form
where $\Theta(\theta-\tau)$ denotes the periodic theta function on the circle, the trilinear product is transformed into a general triple integral of the momentum derivative $P^{\prime}$. The rigorous mathematical foundation for expanding loop functionals in such iterated integrals is provided by Chen’s theorem [35]. Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a strictly complete functional basis. This guaranties that our formal expansion in terms of Magnus forms is exhaustive.
@@ 10.5 Algebraic recurrence for the Momentum Loop Equation
By applying this exact derivative formula, the Makeenko–Migdal loop equation sheds its coordinate-space singularities and collapses into a finite, recursive combinatorial algebra. On the left-hand side, the kinematic operator $Q_{\nu}$ acts via a shuffle product between the 3-index $T^{\alpha\beta\gamma}_{\nu}$ tensor and the lower-order $\mathcal{W}^{(n-3)}$ coefficients. On the right-hand side, the integration over $\theta$ seamlessly sews the split traces back together, yielding an algebraic sum over commutator insertions without requiring cyclic sums or residual functional calculus. These relations are depicted in Fig.13
@@ 10.8 The General Master Field $\hat{X}$ and the Word Space
By the Free Moment-Cumulant Theorem [41], this single, order-independent operator contains an infinite tower of algebraically unconstrained parameters (the free planar cumulants $Q$). Because the free Cuntz algebra contains no finite trace identities, our Master Field Theorem guaranties that the trace of the path-ordered exponential is mathematically unrestricted. It exhaustively spans the continuous functional space of any closed loop without truncation.

### v2

@@ Abstract
We present the exact analytic solution of the Makeenko-Migdal loop equations, thereby solving planar QCD ($N_{c}\rightarrow\infty$) in the continuum limit. The solution is constructed by quantizing internal Majorana fermions (elves) on a rigid Hodge-dual minimal surface. We demonstrate that these worldsheet degrees of freedom provide the exact algebraic mechanism required to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing the planar factorization. In the local limit, the theory reduces to a confining analytic twistor string. By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles. The simplest sector with just one pole yields the exact linear Regge spectrum $m^{2}=\frac{\pi\sigma}{4}(n+1/12)$, matching the experimental Regge trajectories for light mesons ($\pi,K$ trajectories with $n=4J$ and $\rho$ trajectory with n=4J2n=4J-2) within 95% confidence intervals. The empirically correct intercepts $\alpha_{\pi}(0)=-1/48$ and $\alpha_{\rho}(0)=+23/48$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations. Ultimately, this exact geometric quantization explicitly realizes Witten’s Master Field as a critical classical trajectory in twistor space.
@@ 1 Introduction
The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the loop equations. At leading order in the $1/N_{c}$ expansion, the Makeenko-Migdal equations  [^1] Referred to hereafter as the MM equations. While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible. exhibit the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
@@
Remark 1.1.
Equation solving vs. model building The starting point of this work is not a proposed “string model of QCD,” but the planar MM loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD. The aim is constructive: to exhibit explicit variables and an explicit functional measure in which the loop equations admit a finite continuum solution and yield physical observables. The principal ingredients of the present construction are not adjustable assumptions. The rigid Hodge-dual minimal surface enters because strict additivity and an (anti)self-dual area derivative are algebraically required for compatibility with the loop hierarchy, while the Majorana determinant enters because its Pauli-principle sign structure enforces planar factorization.
@@
The culminating result of this paper is the exact reduction of the meson spectroscopy problem to a one-dimensional functional integral over boundary twistor data $(\lambda(z),\mu(z),|z|=1)$. By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory. The path integral exactly localizes at critical resonance energies where the Lefschetz thimbles become degenerate. Because the integration variables are coupled holomorphic twistors, this degeneracy intrinsically possesses a corank of 2, mathematically upgrading generic semiclassical branch cuts into the exact simple poles of hadronic states. Planar QCD thus ceases to be a theory of stochastic quantum fluctuations; the extraction of the exact spectrum reduces to a deterministic, classical generalized eigenvalue problem.
The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral. We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent. Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle. This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$. By analytically continuing the action to the boundary and summing over the exact monodromy phases, we obtain an exact Bohr-Sommerfeld quantization condition. For the simplest sector with just one pole, this yields the spectrum:
$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (1.1)
where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory. Remarkably, the intercepts 1/48,+23/48-1/48,+23/48 emerge directly from the 2D Liouville kinetic anomaly.
In a profound sense, the culminating result of this construction is the explicit geometric realization of Witten’s original "Master Field" hypothesis for large-$N_{c}$ QCD. While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, we discover its exact representation in loop space. The true classical limit of Planar QCD is identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory and accessible via modern geometric quantization.
@@
To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with five specific navigation points that address anticipated physical and computational questions.
@@
Sections 10 to 12 introduce momentum loop space and the twistor string framework. To clarify (ii) the status of Wilson-loop “renormalization,” Sec. 10.1 sharply outlines our perspective: $W[C]$ is kept regularized at a finite Elf mass, physical observables are constructed natively in momentum loop space $W[P]$ (where coordinate-space singularities are integrated out), and only then is the local limit taken. Following this, Sections 11 to  12 execute the central change of variables. Addressing (iii) the parametrization of the quark loop in phase space by boundary values of holomorphic twistors, Sec. 11 details how the linear phase space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings. Section 15 discusses the semiclassical (WKB) limit and the emergence of linear Regge behavior.
Finally, we highlight (iv) the exact analytic extraction of the mass spectrum. Section 16 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory. Section 17 then reframes this localization as an infinite-dimensional realization of Picard-Lefschetz theory, Resurgence, and Exact WKB analysis. We establish a strict structural correspondence between the twistor monodromies of Geometric QCD and the spectral networks of Seiberg-Witten theory, identifying the exact resonance condition as the twistor-string analogue of an Argyres-Douglas point. Section 18 concludes by synthesizing these results into the final geometric Master Field. The appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C and computations of the Minkowski action for our chiral helicoid minimal surface.
Remark 1.2.
Conceptual Demarcation: Bypassing 2D Quantum Gravity
Because the mathematical vocabulary of this paper (strings, twistors, minimal surfaces) intersects with decades of literature on string theory, a reader may instinctively attempt to map this construction onto the conventional quantization of a fluctuating 2D worldsheet. To prevent fundamental conceptual misunderstandings, it is vital to explicitly state that we are not proposing a phenomenological string model, nor are we quantizing a fundamental string in the Polyakov sense.Consequently, the reader will not find integrations over fluctuating 2D bulk metrics, conformal ghost sectors, or integrations over ambiguous Teichmüller moduli spaces. Attempting to evaluate this work by the standard criteria of abstract 2D string modeling is a category error.Our absolute starting point is the rigorously defined, 1-dimensional phase space of Planar QCD: the momentum-loop representation of the MM equations, governed by the exact linear measure in phase space $\int\mathcal{D}C^{\prime}\mathcal{D}P$. The transition to "twistor space" is strictly a local, 1-dimensional algebraic change of variables on the boundary circle. The twistors do not dynamically fluctuate across a 2D bulk; they are localized strictly at the 1D boundary to parametrize the loop velocity, and are analytically continued inward solely as a tool of complex analysis. We are transforming the intractable functional geometry of a 1D gauge-theory phase space into a rigorously solvable problem of 1D complex analysis and linear algebra.
@@ 2.3 The Physical Vacuum and Multi-Instanton Resummation
Before we formulate the dynamics of the Fermi string on the minimal surface, we must clarify the physical origin of this rigid geometric background. The MM (MM) loop equations serve as the exact quantum equations of motion for Planar QCD. Like any equations of motion, they possess an inherent ambiguity corresponding to the choice of the physical vacuum. The mathematical freedom to multiply the perturbative loop functional by a zero mode exactly reflects this physical vacuum ambiguity.
@@ 3.3 The Virasoro Constraint and Uniformization
In the presence of the Virasoro constraint (3.21), the problem becomes a challenging one and has been investigated over the last two centuries by great mathematicians, starting with Riemann and Hilbert. We give a brief summary of modern state of this theory in B.
@@ 6.6 Loop equation for fermion determinants
This is clear from the original representation of $\Delta$ as an exponential of the sum over configurations of independent loops. This planar hierarchical set of loops is shown in Fig. 4, with the red line illustrating the time slice: the number of Elf pairs flying between quarks.
Figure 4: The hierarchical set of loops, some touching, but never intersecting each other nor self-intersecting. The time slice (red line) represents some number of Elf pairs.
@@ 10.2 Momentum loop path integral
Remark 10.1.
Phase Space Normalization and the Departure from the Brownian Gauge.
To appreciate the mathematical control within this phase-space formulation, it is instructive to examine its exact canonical normalization. The functional measure is strictly normalized by the standard symplectic volume element $\mathcal{D}P\mathcal{D}C^{\prime}/(2\pi)$ per continuous degree of freedom. Consequently, in the absence of the interacting Wilson loop (the free-particle limit, $W[C]=1$), integrating over the loop coordinates $C$ natively yields a functional Dirac delta function, strictly enforcing the conservation of local momentum:
$\int d^{4}k\int\frac{\mathcal{D}C^{\prime}\mathcal{D}P}{2\pi}\exp\left(\mathrm{i}\oint(P_{\mu}(\tau)-k_{\mu})\dot{C}_{\mu}d\tau\right)F[P]$
$=\int d^{4}k\int\mathcal{D}P\,\delta[P(\cdot)-k]F[P]=\int d^{4}k\,F[k]$ (10.4)
This demonstrates that the phase-space integral is rigorously well-defined from standard quantum mechanics, unique strictly modulo the infinite 1D diffeomorphism volume $\text{Vol}(\text{Diff}(S^{1}))$ associated with the simultaneous reparametrization of $P(\tau)$ and $C(\tau)$.Conventionally, this $\text{Diff}(S^{1})$ volume is factored out by imposing a proper-time gauge-fixing via an $\int(C^{\prime})^{2}d\tau$ kinetic term. This specific gauge choice inevitably generates the Feynman-Schwinger representation with its standard Brownian (Wiener) measure in loop space—a measure inherently dominated by everywhere-discontinuous fractal paths that artificially generate the notorious UV cusp singularities. In our framework, we recognize that the Brownian measure is just one possible choice of gauge fixing for the 1D diffeomorphisms. We will deliberately abandon it. Instead, as detailed in Section 11, we fix $\text{Diff}(S^{1})$ by imposing the Virasoro constraint $(f^{\prime})^{2}=0$ on the holomorphic twistors parametrizing the loop velocity. This specific gauge choice entirely avoids the fractal artifacts of the Brownian measure and perfectly locks the 1D phase-space kinematics of the quark loop to the 2D internal confining geometry of the minimal surface.
@@ 10.4 Equivalence of loop diffusion operator to the third Magnus form
where $\Theta(\theta-\tau)$ denotes the periodic theta function on the circle, the trilinear product is transformed into a general triple integral of the momentum derivative $P^{\prime}$. The rigorous mathematical foundation for expanding loop functionals in such iterated integrals is provided by Chen’s theorem [3]. Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a strictly complete functional basis. This guarantees that our formal expansion in terms of Magnus forms is exhaustive.
@@ 10.5 Algebraic recurrence for the Momentum Loop Equation
By applying this exact derivative formula, the MM loop equation sheds its coordinate-space singularities and collapses into a finite, recursive combinatorial algebra. On the left-hand side, the kinematic operator $Q_{\nu}$ acts via a shuffle product between the 3-index $T^{\alpha\beta\gamma}_{\nu}$ tensor and the lower-order $\mathcal{W}^{(n-3)}$ coefficients. On the right-hand side, the integration over $\theta$ seamlessly sews the split traces back together, yielding an algebraic sum over commutator insertions without requiring cyclic sums or residual functional calculus. These relations are depicted in Fig.13
@@ 10.8 The General Master Field $\hat{X}$ and the Word Space
By the Free Moment-Cumulant Theorem [48], this single, order-independent operator contains an infinite tower of algebraically unconstrained parameters (the free planar cumulants $Q$). Because the free Cuntz algebra contains no finite trace identities, our Master Field Theorem guarantees that the trace of the path-ordered exponential is mathematically unrestricted. It exhaustively spans the continuous functional space of any closed loop without truncation.
@@ 11 The twistor parametrization of the loop space measure
Remark 11.1.
Measure Transformation via Algebraic Jacobians.
As emphasized in Remark 1.2, we are not evaluating an abstract functional measure over a fluctuating 2D surface. Therefore, the measure we derive here does not encounter the ambiguities of Teichmüller space. Our starting point is the rigorously defined, linear measure of the 1D quark-loop phase space: $\int\mathcal{D}C^{\prime}\mathcal{D}P$. The transition to twistor space is strictly a local (in $\theta$) change of variables, mapping the bounding curve velocity $C^{\prime}(\theta)\propto\text{Im}(z\lambda\sigma\mu)|_{z=e^{i\theta}}$ directly to the boundary twistors $\lambda,\mu$. Because the twistors are localized at the boundary, this transformation involves no functional anomalies. Once the zero modes are fixed, we derive the corresponding twistor measure via standard, explicitly calculable algebraic matrix Jacobians.
@@ 11.2 The residual $U(1)$ gauge symmetry and normalized twistors
11.3 Euclidean Self-Intersection vs. Minkowski Helicity
A common objection to minimal surfaces in the context of QCD strings arises from Euclidean intuition, where a surface of finite area bounded by a loop must satisfy strict embedding theorems (e.g., the Meeks-Yau theorem). In the Euclidean "rubber band" analogy, a surface cannot pass through itself without creating a physical self-intersection, which would typically be singular or unstable.
Euclidean Space: The surface is constrained to a finite sheet. Any attempt to introduce a logarithmic branching point would force the sheet to intersect itself.
Minkowski/Twistor Space: The surface behaves like a helicoid. As the trajectory winds around the twistor pole $\lambda(z)\propto 1/(z-z_{0})$ or $\mu(z)\propto 1/(z-z_{0})$, it moves to a "different sheet" in the complexified coordinate space. This is why the surface can have an infinite area in the Minkowski sense—it is effectively a Riemann surface with an infinite number of sheets: the coordinate map function $f_{\mu}(z)\propto\log(z-z_{0})$.
This multiple-valued coordinate is what allows the poles to exist without the "self-intersection" catastrophe. The surface doesn’t intersect itself for the same reason a spiral staircase doesn’t intersect the floor below it: it has moved in an orthogonal dimension. The "infinite area" is simply the total area of all these sheets, and the monodromy is the "height" gained between each floor of the staircase, which we quantize as:
$\Delta S_{eff}=\int_{sheet_{n}}^{sheet_{n+1}}\mathcal{L},d\tau=2\pi n$ (11.12)
In this complexified geometry, the ’infinite area’ is not a divergence of the physical action, but the sum over the topological winding numbers that localizes the energy poles $E_{n}$ at the discrete values where the monodromy satisfies the phase coherence condition (see the Section 15).